OpenStudy (anonymous):

If g(x) = 3x + 1, evaluate g(-14). a.-41 b.-10 c.-5

OpenStudy (aravindg):

Plug in -14 for x. You will get the answer.

OpenStudy (anonymous):

how about If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(2) + g(3). a.5/3 b.12 c.15

OpenStudy (aravindg):

Did you get the answer for previous question?

OpenStudy (anonymous):

yes -41

OpenStudy (aravindg):

How?

OpenStudy (nurali):

f(x)= x^2 + 1 f(2)= (2)^2+1 f(2)=4+1 f(2)=5 g(x) = 3x + 1 g(3) =3(3)+1 g(3)=9+1 g(3)=10. f(2) + g(3) 5+10 ______________

OpenStudy (anonymous):

f(x) = x 2 + 1 and g(x) = 3x + 1, find f(2) + g(3) This is a 2 part question. So f(2) = (2)^2 +1 and g(3) = 3(3) + 1 so f(2) = 5 and g(3) = 10 Last question: find f(2) + g(3) 5 + 10 = 15

OpenStudy (anonymous):

what about If f(x) = x^2 + 1 and g(x) = 3x + 1, find [f(2)-g(1)] ^2. a.1 b.2 c.9

OpenStudy (anonymous):

Same process as last question. Plug in 2 for f(x), 1 for g(x) f(2) = (2)^2 + 1 = 4 +1 = 5 g(1) = 3(1) + 1 = 3 +1 = 4 Plug in f(2) = 5 and g(1) = 4 to [f(2)-g(1)] ^2. you get [5-4]^2, by using PEMDAS we get [1]^1 = 1 * 1 = "1"